OpenStudy (anonymous):

4r(2) + 7r + 3s(2) 11r(3) + 3s(2) Are these expressions equivalent?

OpenStudy (phi):

does 4r(2) mean \( 4r^2 \) if so, use an ^ to show it is an exponent like this: 4 r^2

OpenStudy (anonymous):

Yes, that's what I meant. I didn't know up to make them like that off my computer. Are they equivalent?

OpenStudy (phi):

they are asking if 4r^2 + 7r is the same as 11 r^3 if you tried with r=2, you would get 4*2^2 + 7*2 =? 11*2^3 or 4*4 + 14 =? 11*8 16+14= 88 30=? 88 ... It did not work. so in general it is not true

OpenStudy (phi):

but you should learn that you cannot combine terms that have different exponents. x^1 + x^2 cannot be combined. on the other hand x^2 + x^2 can be combined (you have 2 x^2)

OpenStudy (anonymous):

Okay. Could you help me with this one? 5mn - 8mn +13mn, simplify.

OpenStudy (anonymous):

would it be 10mn?